Score on last try: 0 of 1 pts. See Details for You can retry this question below Solve -5 +3 + 10 = 7 for I 50 9 X
Answer #1

5 on both * - 5+3 5x+10 = 7 m, -5+ 3 5x+10 + 5 3 J2+10 + 5 = 4 +5 Adding or 3 Jat10 Sides to cancel - 5 = 12 m, by 3 3 Ix+10